The Big Tool Company has budgeted sales of $300,000 with the following budgeted costs:

Direct materials $60,000

Direct manufacturing labor 35,000

Factory overhead

Variable 30,000

Fixed 45,000

Selling and administrative expenses

Variable 20,000

Fixed 25,000


Compute the average markup percentage for setting prices as a percentage of the Variable manufacturing costs.
Select one:
a. 106.8%

b. 140%

c. 39.5%

d. 76.5%
Answer #1

Answer is option b)140%

Sale values =300,000

Total variable manufacturing cost = 60000+35000+30000 = 125000

average markup percentage for setting prices as a percentage of the Variable manufacturing costs = (300000-125000)/125000 =140%